# Do More Suspicious Transaction Reports Lead to More Convictions for Money Laundering?

### v1

@@ 5.2 Fixed Effects Regression Results
F1TF 113 -0.02 27.14*** -0.08 (0.20)
F2TF 113 -0.01 22.49*** -0.08 (0.21) 1.71 (5.29)
F3TF 113 -0.00 15.56*** -0.07 (0.21) 1.95 (6.08) 0.28 (1.89)
F4TF 111 -0.00 14.15*** -0.02 (0.26) -0.36 (3.95) -0.04 (2.12) -1.72* (0.82)
Dependent variable: $\log_{10}(CON_{\left(c,t+1:3\right)})$
F1TF 122 0.10 28.53*** 0.20 (0.19)
F2TF 122 0.09 21.67*** 0.19 (0.19) -1.12 (5.62)
F3TF 122 -0.03 21.18*** -0.03 (0.17) -6.06 (5.59) -4.59** (1.50)
F4TF 116 0.03 21.39*** 0.01 (0.15) -4.47 (6.26) -4.44** (1.62) -1.40 (0.87)
Dependent variable: $\log_{10}(CON_{\left(c,t+1:5\right)})$
F1TF 80 0.08 32.81*** 0.18 (0.10)
F2TF 80 0.11 26.24*** 0.19 (0.10) 3.08 (5.95)
F3TF 80 -0.03 20.45*** 0.09 (0.07) 1.63 (5.87) -2.51 (1.90)
F4TF 80 0.04 22.67*** 0.08 (0.04) 2.80 (5.52) -2.35 (1.77) -2.22 (1.55)
Table 9: Fixed effects regression results with both country- and time-specific effects. Note that standard errors are clustered both at the country and time level. The table contains results for multiple model specifications, F1TF to F4TF, and multiple dependent variables. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level. Note that negative $R^{2}$ (within) values can occur due to the way the statistic is computed; see the linearmodels documentation for details [65].
@@ 6 Conclusion and Discussion
We stress that our analysis focuses on the relationship between the number of STRs and convictions for money laundering in a country. We do not consider the size of convictions (i.e., if they pertain to big or small money laundering schemes or if significant sums are confiscated). We have no doubt that some STRs lead to important convictions and without insight into the context and magnitude of convictions, simple conviction counts might fail to capture important aspects of AML efforts. Furthermore, as one employee of the Danish financial intelligence unit pointed out to us, STRs may be used to secure convictions for other crimes than money laundering such as tax evasion. Our analysis does not capture this important benefit of STRs or how STRs may be used to further how financial intelligence units and reporting entities understand money laundering techniques. After the end of our STR data collection period, the EU introduced its 4th AML directive in an attempt to homogenize AML legislation and promote a risk-based approach. Our data does not allow us to evaluate whether or not this was successful. However, as discussed in Section 2, we note that the volume of STRs in the EU does not appear to have decreased after the 4th directive was introduced. This only emphasizes the importance of “quality vs. quantity considerations” in relation to STRs.

### v2

@@ 5.2 Fixed Effects Regression Results
FETF1 113 -0.02 27.14*** -0.08 (0.20)
FETF2 113 -0.01 22.49*** -0.08 (0.21) 1.71 (5.29)
FETF3 113 -0.00 15.56*** -0.07 (0.21) 1.95 (6.08) 0.28 (1.89)
FETF4 111 -0.00 14.15*** -0.02 (0.26) -0.36 (3.95) -0.04 (2.12) -1.72* (0.82)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:3)}\right)$
FETF1 122 0.10 28.53*** 0.20 (0.19)
FETF2 122 0.09 21.67*** 0.19 (0.19) -1.12 (5.62)
FETF3 122 -0.03 21.18*** -0.03 (0.17) -6.06 (5.59) -4.59** (1.50)
FETF4 116 0.03 21.39*** 0.01 (0.15) -4.47 (6.26) -4.44** (1.62) -1.40 (0.87)
Dependent variable: $\log_{10}\left(CON_{(c,t+1:5)}\right)$
FETF1 80 0.08 32.81*** 0.18 (0.10)
FETF2 80 0.11 26.24*** 0.19 (0.10) 3.08 (5.95)
FETF3 80 -0.03 20.45*** 0.09 (0.07) 1.63 (5.87) -2.51 (1.90)
FETF4 80 0.04 22.67*** 0.08 (0.04) 2.80 (5.52) -2.35 (1.77) -2.22 (1.55)
Table 9: Fixed effects regression results with both country- and time-specific effects. Note that standard errors are clustered both at the country and time level. The table contains results for multiple model specifications, named FETF1 to FETF4, and multiple dependent variables. Columns include model ID, the number of observations $n$ used to fit each model, goodness of fit $R^{2}$ (within), F-statistic for poolability, and coefficient estimates (standard errors in parentheses). We use * to denote significance at the $0.05$ $p$-level; ** at the $0.01$ $p$-level, and *** at the $0.001$ $p$-level. Note that negative $R^{2}$ (within) values can occur due to the way the statistic is computed; see the linearmodels documentation for details [65].
@@ 6 Conclusion and Discussion
We stress that our analysis focuses on the relationship between the number of STRs and convictions for money laundering in a country. We do not consider the size of convictions (i.e., if they pertain to big or small money laundering schemes or if significant sums are confiscated). We have no doubt that some STRs lead to important convictions and without insight into the context and magnitude of convictions, simple conviction counts might fail to capture important aspects of AML efforts. Our analysis also does not capture how STRs might help improve financial intelligence units’ and reporting entities’ understanding of money laundering. Furthermore, as one employee of the Danish Financial Intelligence Unit pointed out to us, STRs may be used to secure convictions for other crimes than money laundering (for example tax evasion). Our analysis does not addressed or capture this. However, in Appendix C, we consider three types of predicate crimes to money laundering: corruption, drug trafficking, and fraud. While the appendix does not rule out the possibility of a positive relationship between the number of STRs and convictions for predicate crimes, it also does not find clear, robust, and consistent evidence of such a relationship.
After the end of our STR data collection period, the EU introduced its 4th AML directive in an attempt to homogenize AML legislation and promote a risk-based approach. Our data does not allow us to evaluate whether or not this was successful. However, as discussed in Section 2, we note that the volume of STRs in the EU does not appear to have decreased after the 4th directive was introduced. We believe that this only emphasizes the importance of “quality vs. quantity considerations” in relation to STRs.